John needs 40 Eternal Light flashlights. Each Eternal Light flashlight requires 3 bulbs, and each bulb requires 3 filaments. Note that the bulbs are non-sealed and require the filaments to be inserted inside of them as a sub-component.

1. How many completed bulbs will he need, if he has 90 in inventory

A. 30

B. 120

C. 240

2. How many filaments will he need, if he has 90 completed bulbs in inventory and 20 filaments in inventory?

A. 110

B. 90

C. 70

     

3. If John is looking at a report which shows that his filaments will be late in arriving, he is looking at a

A. Action Notice - an Output of the MRP

B. Inventory Record - an Input to the MRP

C. Capacity Report - an Output of the MRP

4. If John in the Accounting Department wants to determine product costing, and easily share that with the Finance Department he would use

A. an MRP

B. an ERP

C. an MRI

  

5. If John knows he will need 40 flashlights by December 8th and documents it,

A. that would constitute a Master Production Schedule, an Input to the MRP

B. that would be a Performance Report, an Output of the MRP

C. that would be described as a Bill of Materials, an Input to the MRP
Answer #1

Each flashlight needs 3 bulbs

Net Bulbs required = Bulbs required - Inventory = 40*3 -90 = 30 bulbs

Each bulb needs 3 filaments.

Net Filaments required = 30 *3 - On-hand filaments = 90-20 = 70 filaments

3. John is looking at an action notice that informs if a part is getting delayed or due dates may get pushed out.

4. John would use an ERP. MRP is used for material planning and contains no financial information.

5. Master production schedule gives a schedule of production required to meet demand. Answer = Option a.